OpenStudy (anonymous):

Solve, assuming a linear equation fits the situation. The value of a graphing calculator is $225. After 2 years, the value of this calculator is $160. Find the value of the calculator after 5 years. (HINT: Let the time now be 0, the time in two years be 2, and the time in five years be 5.)

OpenStudy (anonymous):

to get the linear equation we have to get the slope and the starting y value (whats it called in english?) starting y is obviously 225 slope is (160-225)/2 = -32.5 so the equation would be: f(x)=-32.5x+225 with x being years and f(x) being the value after x years. so after 5 years its value would be: f(5)=-32.5*5+225=62.5 so its value after 5 years would be 62.5$

OpenStudy (anonymous):

62.5 is rong ... i typed it in and it said it was rong

OpenStudy (anonymous):

62.5 is rong ...i typed it on and it said it was rong

OpenStudy (anonymous):

Try typing in 162.5.

OpenStudy (anonymous):

call me over delaine i'm bored.

OpenStudy (anonymous):

are you sure? in 2 years it lost $225-$160=$65 value. so in 1 year it loses $65/2 = $32.5 (assuming a linear equation). so in 5 years it loses $32.5*5=$162.5. so after 5 years it has $225-$162.5=$62.5

OpenStudy (anonymous):

Make x the price of the calculator at year 0 x=225 Year 2 equation. Y represents the year x=160+2y Solve simultanoeusly 225-160+2y=0 62=2y 32.5*5= 162.5....... Yeah the answer should be $62.5